OpenStudy (anonymous):

simplify the expression 2 over 2+square root of 3

jhonyy9 (jhonyy9):

do you know what mean the conjugate of 2+sqrt3 ?

OpenStudy (anonymous):

no

jhonyy9 (jhonyy9):

but i think that you have learned

OpenStudy (anonymous):

could you explain?

jhonyy9 (jhonyy9):

so do you know formula a^2 -b^2 = ?

jhonyy9 (jhonyy9):

factorized

OpenStudy (anonymous):

yeh

jhonyy9 (jhonyy9):

so than write it here now please

jhonyy9 (jhonyy9):

courage i will help you

jhonyy9 (jhonyy9):

so your exercise is 2 ---------- = 2 + sqrt 3

OpenStudy (anonymous):

i dont understand like can you work it out?

jhonyy9 (jhonyy9):

so this i wann you learne it now

OpenStudy (anonymous):

yeah I learn when I see things worked out

jhonyy9 (jhonyy9):

so there is this fraction with radical on determinant and when you need to simplify it this meaning that you need meking calculi so than the radical will be eliminated from determinant this is ok ?

jhonyy9 (jhonyy9):

so there is on determinant 2 +sqrt3 when i have asked you that you know it sure formula a^2 -b^2 = (a-b)(a+b) here is factorized

jhonyy9 (jhonyy9):

than this meaning that when you multiplie the (a-b) by conjugate what is (a+b) so than you will get like result a different of two square

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

so in case of your exercise this will be so there is on determinante 2+sqrt3 so 2+sqrt3 has a conjugate of 2-sqrt3 becasue (2+srt3)(2-sqrt3)=4-3 =1 than your exercise will be 2 2(2 -sqrt3) 4 - 2sqrt3 4 -2sqrt3 ------------ = ----------------- = --------------- = -------------- 2 +sqrt3 (2+sqrt3)(2-sqrt3) 4 - 3 1 = 4 -2sqrt3 do you understand it ?